OpenStudy (anonymous):

Solve the system using substitution and describe your steps x = y - 3 2x + y = 12

Vocaloid (vocaloid):

the first equation (x = y - 3) tells us that we can replace x with y - 3 whenever we see it, right? look at the second equation, replace x with (y-3) and solve for y then go back and solve for x using the first equation

OpenStudy (anonymous):

i need more help

Vocaloid (vocaloid):



Vocaloid (vocaloid):

does this make sense? see how we replace "x" with (y-3)?

OpenStudy (anonymous):

is that the final answer?

Vocaloid (vocaloid):

no, but do you understand so far?

OpenStudy (anonymous):

yes

Vocaloid (vocaloid):

ok, so we have 2(y-3) + y = 12, so what do you get when you distribute the 2?

OpenStudy (anonymous):

where do u distribute it to

OpenStudy (anonymous):

this is a new concept

Vocaloid (vocaloid):



OpenStudy (anonymous):

2(y-3) = 2*y-6

Vocaloid (vocaloid):

excellent, so far we have 2y - 6 + y = 12, so combine like terms to find y

OpenStudy (anonymous):

3y-6=12

Vocaloid (vocaloid):

getting there, now add 6 to both sides and divide by 3 to find y

OpenStudy (anonymous):

6

Vocaloid (vocaloid):

great, so we know that y = 6, but we're still not done yet because we need to find x from the first equation, we know that x = y - 3, and since y = 6, what is x?

OpenStudy (anonymous):

3?

Vocaloid (vocaloid):

good! x = 3 and y = 6 generally, the solution to a system of equations is written as (x,y), so since x = 3 and y = 6, our solution is (3,6) that's our final answer

OpenStudy (anonymous):

thank you